OpenStudy (anonymous):

Write the complex number in the form a + bi. square root of six(cos 315° + i sin 315°)

OpenStudy (anonymous):

I dont know how to solve this at all... My choices: square root of three over two minus square root of thee over two times i square root of six minus square root of sixi square root of three minus square root of threei square root of six over two minus square root of six over two times i

Nnesha (nnesha):

look at the unit circle cos 315 = what ? remember (x,y) sin =y-coordinate and cos = x-coordinate

OpenStudy (anonymous):

cos315 is square root(2)/2

OpenStudy (anonymous):

so \[\frac{ \sqrt{2} }{ 2 }\] is my x coordinate?

Nnesha (nnesha):



OpenStudy (anonymous):

Yes, I was just looking at that

Nnesha (nnesha):

\[\huge\rm \sqrt{6}(\cos 315° + i \sin 315°)\] so cos 315 =sqrt{2} over 2 replace cos 315 by that

OpenStudy (anonymous):

So i would replace isin315 with \[\frac{ -\sqrt{2} }{ 2 }\]

OpenStudy (loser66):

@Nnesha Good job!! open my eye!! :)

Nnesha (nnesha):

\[\huge\rm \sqrt{6}(\color{reD}{cos 315°} + i \sin 315°)\] \[\sqrt{6}(\frac{ \sqrt{2} }{ 2 }+i (-\frac{ \sqrt{2} }{ 2 }))\] ust sin315 not the i

Nnesha (nnesha):

just**

OpenStudy (anonymous):

Alright, then do I distribute the \[\sqrt{6}\]?

Nnesha (nnesha):

thanks o^_^o @66

Nnesha (nnesha):

first distribute by i i times -sqrt{2} over 2

Nnesha (nnesha):

\[\sqrt{6}(\frac{ \sqrt{2} }{ 2} - \frac{ \sqrt{2} }{ 2}i)\]

OpenStudy (anonymous):

Okay, after distributing i, do I distribute the square root of 6?

Nnesha (nnesha):

well i wouldn't do that just take out the common factor

Nnesha (nnesha):

\[\sqrt{6}\color{ReD}{(\frac{ \sqrt{2} }{ 2} - \frac{ \sqrt{2} }{ 2}i)}\] what is common factor in the parentheses

OpenStudy (anonymous):

Hmm... then my end result is \[\sqrt{3}(-i+1)\]

Nnesha (nnesha):

u sure or just guessing ? ;P

OpenStudy (anonymous):

100% sure. I just calculated it. If I distribute the square root of 3, then I get what appears to be my third choice. Am I correct?

Nnesha (nnesha):

alright yes that's correct there are square roots that's why i don't like distributing by sqrt{6} you can take out the common factor which is sqrt{2}/2

OpenStudy (anonymous):

Oh, I see. @Nnesha

Nnesha (nnesha):

\[\sqrt{6} \times \frac{ \sqrt{2} }{ 2 }(1-i)\] \[\frac{ \sqrt{12} }{ 2 }(1-i)\]multply sqrt{6} times sqrt{2} now factor 12 \[\frac{ \sqrt{4 \times 3} }{ 2}(1-i)\] take the square root of 2 you will get the same answer

Nnesha (nnesha):

i found it easy but you can apply any method as lng s you get the same answer